# Affine Frequency Division Multiplexing (AFDM) for 6G: Properties, Features, and Challenges

### v1

@@ III-H Inherent Physical-Layer Security
AFDM provides inherent physical-layer security through configurable the chirp parameters ($c_{1}$, $c_{2}$) and transform-domain methods such as chirp permutation [11], without degrading performance.
Legitimate users can treat these configurations as shared secret keys, while unauthorized receivers face infeasible brute-force search over a continuous parameter space, shown to be infeasible even under quantum-accelerated attackes. By contrast, OFDM employs fixed DFT-based subcarriers with no intrinsic secrecy, and while OTFS offers limited structural confidentiality via discrete grid size options, but the finite configuration space is constrained by channel diversity requirements. AFDM, with its broader and continuous parameter space, enables significantly stronger waveform-level confidentiality.
@@ H. Testbeds, Trials, and Reference Designs
The preceding sections have provided a comprehensive overview of the AFDM waveform, outlining its key signal-theoretic features and practical advantages across multiple domains, including channel resilience, diversity, spectral properties, and computational efficiency. Building upon this foundation, this section explains how these technical features align with some emerging and current 6G requirements and design principles outlined by standardization bodies and academia, such as 3GPP [2], International Telecommunication Union - Radiocommunication Sector (ITU-R) [3], and Institute of Electrical and Electronics Engineers (IEEE).
@@ IV-D Ultra-Low Latency: Implementation Complexity
The practical adoption of any new waveform hinges on its ability to support real-time processing under stringent latency and complexity constraints, especially in 6G systems that demand extremely low latency and high throughput. Although AFDM relies on the advanced DAFT, it was shown - as discussed in previous sections - to be implemented efficiently using structured fast transform algorithms, remaining tractable for even larger systems increasingly expected in 6G networks.
@@ IV-F Integration with Emerging Use Cases
The adoption of any new waveform in 6G depends on its ability to reliably and efficiently support emerging use cases and service requirements. AFDM demonstrates strong potential in this regard, owing to its inherent flexibility, delay-Doppler robustness, and processing efficiency, well-aligned with key 6G paradigms.
For instance, AFDM is particularly well-suited for high-mobility scenarios such as V2X communications and SAGIN architectures involving low Earth orbit (LEO) constellations and unmanned aerial vehicles (UAVs), due to its resilience to Doppler shifts and delay spreads, supporting the vision of key verticals like automotive and future vehicular services [16].
@@ IV-G Robustness to Practical Impairments
Robustness to real-world impairments remains a key requirement for waveform standardization, particularly for deployment in diverse hardware environments and under non-ideal channel conditions. Such impairments include carrier frequency offset (CFO), phase noise, amplifier nonlinearities, and quantization effects, all of which are known to degrade multicarrier waveforms like OFDM.
As evaluated, the AFDM exhibits natural resilience to these challenges, where the spreading in the DAFT domain distributes energy across time and frequency, mitigating sensitivity to CFO and providing improved phase noise tolerance. These properties position AFDM as a robust and scalable candidate for practical 6G scenarios where performance must be preserved despite front-end imperfections.

### v3

@@ III-H Inherent Physical-Layer Security
AFDM provides inherent physical-layer security through configurable the chirp parameters ($c_{1}$, $c_{2}$) and transform-domain methods such as chirp permutation without degrading performance. Legitimate users can treat these configurations as shared secret keys, while unauthorized receivers face infeasible brute-force search over a continuous parameter space, shown to be infeasible even under quantum-accelerated attacks. By contrast, OFDM employs fixed DFT-based subcarriers with no intrinsic secrecy, and while OTFS offers limited structural confidentiality via discrete grid size options, but the finite configuration space is constrained by channel diversity requirements. AFDM, with its broader and continuous parameter space, enables significantly stronger confidentiality.
@@ H. Testbeds, Trials, and Reference Designs
The preceding sections have provided a comprehensive overview of the AFDM waveform, outlining its key signal-theoretic features and practical advantages across multiple domains, including channel resilience, diversity, spectral properties, and computational efficiency. Building upon this foundation, this section explains how these technical features align with some emerging and current 6G requirements and design principles outlined by standardization bodies and academia, such as 3GPP [2, 13], International Telecommunication Union - Radiocommunication Sector (ITU-R) [3], and Institute of Electrical and Electronics Engineers (IEEE).
@@ IV-D Ultra-Low Latency: Implementation Complexity
The practical adoption of any new waveform hinges on its ability to support real-time processing under stringent latency and complexity constraints, especially in 6G systems that demand extremely low latency and high throughput.
Although AFDM relies on the advanced DAFT, it was shown - as discussed in previous sections - to be implemented efficiently using structured fast transform algorithms, remaining tractable for even larger systems increasingly expected in 6G networks.
@@ IV-F Integration with Emerging Use Cases
The adoption of any new waveform in 6G depends on its ability to reliably and efficiently support emerging use cases and service requirements. AFDM demonstrates strong potential in this regard – especially under the ISAC framework – owing to its inherent flexibility, delay-Doppler robustness, and processing efficiency, well-aligned with key 6G paradigms.
For instance, AFDM is particularly well-suited for high-mobility scenarios such as V2X communications and SAGIN architectures involving low Earth orbit (LEO) constellations and unmanned aerial vehicles (UAVs), due to its resilience to Doppler shifts and delay spreads, supporting the vision of key verticals like automotive and future vehicular services.
@@ IV-G Robustness to Practical Impairments
Robustness to real-world impairments remains a key requirement for waveform standardization, particularly for deployment in diverse hardware environments and under non-ideal channel conditions. Such impairments include carrier frequency offset (CFO), non-white noise, amplifier nonlinearities, and quantization effects, all of which are known to degrade multicarrier waveforms like OFDM.
As evaluated, the AFDM exhibits natural resilience to these challenges, where the spreading in the DAFT domain distributes energy across time and frequency, mitigating sensitivity to CFO and providing enhanced robustness to colored noise. These properties position AFDM as a robust and scalable candidate for practical 6G scenarios where performance must be preserved despite front-end imperfections.